OpenStudy (anonymous):

the difference of a number x and -37 is 19 what is the number?

OpenStudy (redplanet2001):

-18

OpenStudy (xguardians):

X = -56 or -18

OpenStudy (anonymous):

x-(-37) = 19

OpenStudy (redplanet2001):

@xGuardians because of the way the question is worded, listing X first it would be -18

OpenStudy (anonymous):

same thing as x+37 = 19

OpenStudy (anonymous):

so 19-37 = -18

OpenStudy (anonymous):

x = -18

OpenStudy (xguardians):

@redplanet2001 Thanks for catching my mistake!

OpenStudy (anonymous):

shay17 your really good lol